### v1

# Mitigating Malicious Attacks in Federated Learning via Confidence-aware Defense

Abstract
Federated Learning (FL) is an emerging distributed machine learning diagram that allows multiple clients to collaboratively train a global model without sharing private local data. However, FL systems are vulnerable to attacks from malicious clients, who can degrade the global model’s performance through data poisoning and model poisoning. Existing defense methods typically focus on a single type of attack, such as Byzantine attacks or backdoor attacks, and are often ineffective against potential data poisoning attacks like label flipping and label shuffling. Additionally, these methods often lack accuracy and robustness in detecting and handling malicious updates. To address these issues, we propose a novel method based on model confidence scores, which evaluates the uncertainty of client model updates to detect and defend against malicious clients. Our approach is comprehensively effective for both model poisoning and data poisoning attacks, and is capable of accurately identifying and mitigating potential malicious updates from being aggregated. Experimental results demonstrate that our method significantly improves the robustness of FL systems against various types of attacks, also achieving higher model accuracy and stability across various scenarios.
Introduction
Federated Learning (FL) is a distributed machine learning paradigm that allows multiple clients to collaboratively train a global model without sharing their local data. This approach offers significant privacy protection, therefore making it highly attractive for applications in healthcare, finance, and mobile device intelligence (Zhang et al. 2021; Rodríguez-Barroso et al. 2023). However, the decentralized nature of FL also introduces new security challenges: Malicious clients can degrade the global model’s performance through data poisoning and model poisoning attacks, leading to inaccurate predictions and potentially high consequences (Fang and Chen 2023; Kumar, Mohan, and Cenkeramaddi 2023). Effectively detecting and defending against malicious clients is crucial for ensuring the security and robustness of FL systems.
Malicious client attacks can be broadly categorized into model poisoning and data poisoning attacks. Model poisoning attacks involve directly manipulating the model updates sent by the clients to the central server. These malicious clients can alter their model parameters or gradients before sending them to the server, with the intent of corrupting the global model’s performance. Such techniques include Byzantine attacks, where the adversary sends arbitrary or random updates, and backdoor attacks, where the adversary subtly poisons the model to misbehave on specific inputs while maintaining overall performance on regular data (Fang et al. 2020; Bhagoji et al. 2019). On the other hand, data poisoning attacks occur when malicious clients intentionally corrupt their local training data to mislead the model during the training process. Such as by label flipping, where the labels of the training data are intentionally altered, and label shuffling, where the data-label pairs are randomly reassigned (Tolpegin et al. 2020; Xie et al. 2020). Both types of attacks pose significant threats to the integrity and robustness of FL systems, and bring the high demand for effective detection and defense mechanisms to ensure reliable model training.
Existing defense mechanisms primarily focus on addressing specific types of attacks. Byzantine-robust methods like Krum (Blanchard et al. 2017) and Trimmed Mean (Yin et al. 2018) aim to mitigate the impact of outliers among client updates but often fall short when dealing with data poisoning attacks. However, these methods generally assume that the majority of clients are benign, which may not hold true in practical scenarios where the proportion of malicious clients is significant (Fang et al. 2020). Byzantine-robust methods aim to reduce the impact of malicious clients on the global model through various techniques. For example, Krum (Blanchard et al. 2017) selects the update closest to the majority of client updates to defend against malicious attacks. However, this approach is insufficient for complex data poisoning attacks because it cannot comprehensively detect and handle all types of malicious updates (Fang et al. 2020). Existing methods attempt to defend against data poisoning attacks by detecting and correcting anomalous data. For instance, Sun et al. (Sun et al. 2019) proposed mitigating backdoor attacks through clipping and noise addition. However, these methods overlook the uncertainty in model updates and fail to address various types of data poisoning attacks comprehensively.
Due to the complexity of attacks and the inherent uncertainty in FL systems, it is insufficient to address only one type of attack. FL systems face a variety of threats, from data poisoning to model poisoning, each presents unique challenges. This diversity necessitates a versatile defense method that is applicable to any type and severity of attack, whether mild, moderate, or severe. A robust and comprehensive approach is urgently needed to ensure the integrity and reliability of the global model across all scenarios. To meet this demand, we propose a novel defense mechanism based on the client confidence score on its learned local model, which refers to the confidence of the model in its predictions. and by analyzing these scores, we can identify updates that exhibit abnormal levels of uncertainty due to malicious actions. By evaluating the uncertainty of client model updates, our approach can detect and defend against various malicious client attacks, under different attack levels. Specifically, our approach includes the following steps: 1. Collecting Confidence Scores of each Client Update: During each training round, we collect the confidence scores of model updates from each client. 2. Establishing Uncertainty Boundaries: Based on the collected confidence scores, we assess the uncertainty of each client model update and set the boundaries as a reference. 3. Detecting and Handling Malicious Updates: Based on the boundaries, we identify updates with high uncertainty and take appropriate actions, such as discarding or adjusting these updates. Our key contributions are as follows:
  • We introduce a simple yet effective confidence-based malicious client detection method that accurately identifies if any attack happens during local model training. Our approach is versatile and applicable to various types of attacks under different attack intensities, including model poisoning or data poisoning attacks.
  • By detecting malicious attacks on client model updates through confidence scores, our method significantly improves the robustness of FL systems against attacks of different intensities. Whether facing mild, moderate, or severe malicious activities, our approach ensures higher model accuracy and stability.
  • We validate our method through extensive experiments on multiple datasets and demonstrate its effectiveness in enhancing the security and performance of FL models against a wide range of attacks. The results show that our confidence-based defense mechanism outperforms existing methods in terms of robustness and accuracy, in detecting and mitigating malicious updates
Related Work
Federated Learning (FL)
FL is a decentralized machine learning approach that enables multiple clients to collaboratively train a shared global model while keeping their local data private. This diagram is particularly vital in scenarios where data privacy is critical, such as healthcare, finance, and mobile applications (Rahman et al. 2023). In FL, the learning process is divided into two main stages: local training and global aggregation. During the local training phase, each client trains a local model using its private dataset. This involves updating the model parameters based on the local data and ensuring that the raw data never leaves the client’s device. Once the local training is complete, the clients send their updated model parameters to a central server for global aggregation, in which involves the central server aggregates the received model updates from all participating clients to form a new global model. This aggregation can be done by various approaches, with the most common being the Federated Averaging (FedAvg) algorithm (McMahan et al. 2017). The FedAvg algorithm computes the weighted average of the model parameters from all clients to form a global model. Despite the advantages of FL in preserving data privacy, the decentralized nature of this learning paradigm makes it vulnerable to various types of attacks. Given the susceptibility of FL to such attacks, it is crucial to develop efficient and robust defense mechanisms that can identify and mitigate the impact of malicious activities. This is what we bring in the paper.
Malicious Client Attacks
FL models are susceptible to various types of attacks from malicious clients, which can be broadly categorized into data poisoning and model poisoning.
Data Poisoning Attacks: In data poisoning attacks, malicious clients intentionally manipulate their local training data to degrade the performance of the global model. Common strategies include label flipping, where the labels of training samples are deliberately swapped, and label shuffling, where the labels of training samples are randomly reassigned. These attacks aim to bias the model’s training process, and further lead to inaccurate or harmful predictions. Rahman et al. (Rahman et al. 2023) highlights how local model poisoning through data manipulation can significantly impact the robustness of federated learning systems. Similarly, Smith et al. (Smith and Brown 2023) discuss defense mechanisms against clean-label poisoning attacks, where the manipulated data appears legitimate but is designed to subvert the model’s performance.
Model Poisoning Attacks: In contrast to data poisoning attacks, model poisoning attacks refer to malicious clients directly sending manipulated model updates to the central server during the global aggregation phase. These attacks are conducted by altering the model parameters in a way that maximizes the negative impact on the global model. Zhang et al. (Zhang and Chen 2023) demonstrated how to backdoor attacks could be performed by injecting specific patterns into the model updates and enabling attackers to trigger incorrect behaviors under certain conditions. Lee et al. (Lee and Park 2023) analyzed the adversarial impact of model attacks on federated learning models. Additionally, Byzantine attacks, which involve arbitrary or adversarial behavior by clients, can also be classified under model poisoning when clients send random or adversarial updates to the central server. Recent Byzantine-robust methods like Krum++ (Wang and Zhou 2023) and Enhanced Trimmed Mean (Liu and Zhao 2023) aim to filter out these abnormal updates but can struggle when the attacks are sophisticated.
Defense Against Malicious Attacks
To detect and mitigate the aforementioned attacks, numerous works have been proposed (Fang and Chen 2023; Yan et al. 2023; Zhang and et al. 2022). DeFL (Yan et al. 2023) was proposed as a defense mechanism that secures critical learning periods during the training process to reduce the impact of model poisoning attacks. Zhang et al. (Zhang and et al. 2022) introduced FLDetector to integrate statistical analysis, anomaly detection, and machine learning methods to enhance the accuracy of detecting poisoned updates. This multi-faceted strategy improves the reliability and robustness of FL systems against various types of attacks. Despite the improved accuracy on malicious client detection, these existing methods are often tailored to specific types of attacks, either model or data poisoning, and they typically rely on the fragile assumption that the attack intensity is mild (less than 50%), which may not hold in a practical FL system. To resolve these limitations, in this paper, we propose the CAD model to address both types of attacks and varying intensities.
Model Confidence and Uncertainty
Confidence scores measure how certain a model is about its predictions, with higher scores indicating greater certainty and lower scores reflect higher uncertainty. It provides insights into the model’s decision-making process by identifying the predictions that the model is less sure about (Gal and Ghahramani 2016; Lakshminarayanan, Pritzel, and Blundell 2017). Super Loss, a confidence-aware loss designed for curriculum learning, incorporates confidence scores to determine the reliability of each training instance (Castells, Weinzaepfel, and Revaud 2020). In this work, motivated by the observation that malicious attack will disrupt the local model’s convergence against the original global initialization, and making the learned model less confident in the output, we propose to use per-client confidence accumulation as the index to determine the quality of the learned model.
Methodology
Problem Formulation
We consider an FL system with $n$ clients, each with its local dataset $\mathcal{D}_{i}$. The goal is to train a global model $\theta$ by aggregating the local updates $\theta_{i}$ from client $\{i\}_{i=1}^{n}$. Malicious clients can perform two types of attacks: Data Poisoning Attacks: These include label flipping and label shuffling, where the local dataset is intentionally corrupted to mislead the global model. Model Poisoning Attacks: Here, the model updates sent by the clients are manipulated to degrade the global model’s performance. To mitigate these attacks, we propose a Confidence-Aware Defense (CAD) mechanism based on the per-client confidence accumulation ($\{\sigma_{i}\}_{i=1}^{n}$), which reflects the uncertainty of the model’s predictions and helps in identifying malicious updates. The notations are summarized in Table 1.
Table 1: Table of notations in this paper.
Notation Description
$n$ Number of clients
$\mathcal{D}_{i}$ Local dataset of client $i$
$\theta$ Global model parameters
$\theta_{i}$ Local model parameters of client $i$
$\mathcal{L}_{i}^{j}$ Individual loss for sample $j$ in $\mathcal{D}_{i}$
$\sigma_{i}$ Confidence score for client $i$
$S$ Set of confidence scores from all clients
$\mu_{\text{lower}},\mu_{\text{upper}}$ Centroids from k-means clustering
$\mathcal{H}$ Set of honest clients
$\mathcal{M}$ Set of malicious clients
$\mathcal{L}_{\mathcal{H}}$ Data lengths of honest clients
$W_{\text{orig}}$ Original weights based on data lengths
$R$ Re-weight regularization factors
$r_{\text{norm},i}$ Normalized regularization factor
$w_{\text{final},i}$ Final weight for client $i$
W Lambert W function
Preliminary: Super Loss
Super Loss (Castells, Weinzaepfel, and Revaud 2020) introduces a regularization to adjust the loss dynamically based on the confidence score (${\sigma}^{j}$) for sample $j$ in the training set $\mathcal{D}$. The Super Loss is defined as:
$\mathcal{L}_{\text{super}}=\frac{1}{N}\sum_{j=1}^{N}\left(\frac{1}{{\sigma^{j}}^{2}}\mathcal{L}_{j}+\log\sigma_{j}\right),$ (1)
where $\mathcal{L}_{i}$ is the task loss for sample $i$, and ${\sigma}^{j}$ is the confidence score, and $N$ is the number of samples. The confidence score ${\sigma}^{j}$ is obtained by the following closed-form solution:
${\sigma}^{j}=\exp\left(-\text{W}\left(\frac{1}{2}\max\left(-\frac{2}{e},\frac{\mathcal{L}_{i}-\log(C)}{\lambda}\right)\right)\right),$ (2)
where $\log(C)$ is the constant, and $C$ is set to the number of classes, and $\lambda$ is the regularization parameter. Note that we use this formulation solely to estimate the confidence score but not as a regularizer, as in the original design of Super Loss. The rationale is that our goal is to make the CAD model versatile and applicable to various tasks, and to explicitly demonstrate that the effectiveness stems from the detection algorithm rather than modifications to the loss function.
Confidence-Driven Detection
Client Confidence Estimation
For each client, the confidence score is estimated based on their local dataset $\mathcal{D}_{i}$. Referring to Eq. (2), we calculate the accumulated confidence score for all samples in the dataset and then compute an average to obtain the per-client confidence score $\sigma_{i}$. The accumulated confidence score for client $i$ is the sum of the confidence scores for all samples in their dataset:
$\Sigma_{i}=\sum_{j=1}^{N_{i}}\sigma^{j},$ (3)
where $N_{i}$ is the number of samples in $\mathcal{D}_{i}$. Finally, the per-client confidence score $\sigma_{i}$ is obtained by averaging the accumulated confidence scores:
$\sigma_{i}=\frac{\Sigma_{i}}{N_{i}}.$ (4)
This per-client confidence score $\sigma_{i}$ reflects the overall uncertainty of the model’s predictions for client $i$. Once the per-client confidence scores $\sigma$ have been computed, they will be normalized to ensure comparability across different clients by min-max scaling, which transforms the confidence scores into a uniform range [0, 1]. Let $S=\{\sigma_{i}\}_{i=1}^{n}$ represent the set of confidence scores from $n$ clients. The normalized confidence scores $S_{\text{norm}}$ are formulated as:
$S_{\text{norm}}=\frac{S-\min(S)}{\max(S)-\min(S)}.$ (5)
Clustering Based Detection
After normalizing the per-client confidence scores, the next step is to cluster the clients into groups of honest and potentially malicious clients. The rationale for clustering is based on the assumption that clients can be divided into two groups: honest clients with higher confidence scores and potentially malicious clients with lower confidence scores. To achieve this, we apply k-means clustering to the normalized confidence scores $S_{\text{norm}}$. By setting the number of clusters $k$ to 2, we obtain two centroids to establish the lower bound and upper bound:
$\{\mu_{\text{lower}},\mu_{\text{upper}}\}=\text{K-means}(S_{\text{norm}},2).$ (6)
Once the centroids are obtained, the next step is to classify the clients as honest or malicious based on their proximity to the cluster centroids. Specifically, a client $i$ is deemed honest if its normalized confidence score $\sigma_{i}$ is closer to the centroid $\mu_{\text{upper}}$ (representing higher confidence) than to the centroid $\mu_{\text{lower}}$ (representing lower confidence). Conversely, a client is classified as malicious if the opposite holds true:
$\mathcal{H}$ =$=\{i\mid|\sigma_{i}-\mu_{\text{upper}}|<|\sigma_{i}-\mu_{\text{lower}}|\},$ (7)
$\mathcal{M}$ =$=\{i\mid|\sigma_{i}-\mu_{\text{lower}}|\leq|\sigma_{i}-\mu_{\text{upper}}|\}.$
By classifying clients in this manner, we can focus on aggregating updates from honest clients, thereby enhancing the robustness and security of the federated learning system.
Re-weighted Aggregation
Let $\Theta_{\mathcal{H}}$ represent the set of trained models and $\mathcal{L}_{\mathcal{H}}$ represent the set of data lengths for the honest clients. These are defined as:
$\Theta_{\mathcal{H}}=\{\theta_{i}\mid i\in\mathcal{H}\},$ (8)
=$\mathcal{L}_{\mathcal{H}}=\{\ell_{i}\mid i\in\mathcal{H}\},$ (9)
by which we isolate the local models of clients that are considered trustworthy, so as to ensure that only reliable information is used in the aggregation process. Next, we define the original weights $w_{\text{orig},i}$ based on the data lengths of the honest clients. The original weight for each client is calculated as the ratio of the client’s data length to the total data length of all honest clients:
$W_{\text{orig}}=\{w_{\text{orig},i}\}_{i\in\mathcal{H}}=\{\frac{\ell_{i}}{\sum_{j\in\mathcal{H}}\ell_{j}}\}_{i\in\mathcal{H}}.$ (10)
Subsequently, to adaptively optimize the weight based on the trustworthy of each honest client, a re-weight regularization factor $r_{i}$ is then calculated for each honest client based on their normalized confidence scores. The set of re-weight regularization factors for all honest clients is denoted as:
$R=\{\sigma_{i}\mid i\in\mathcal{H}\}$ (11)
The factors are then normalized:
$r_{\text{norm},i}=\frac{\sigma_{i}}{\sum_{j\in\mathcal{H}}\sigma_{j}}.$ (12)
by using which we can ensure that the aggregation process give appropriate importance to each client model. After calculating the original weights and the re-weight regularization factors, we combine them to obtain the final weights. The final weights are calculated as:
$w_{\text{final},i}=r_{\text{norm},i}\cdot w_{\text{orig},i}.$ (13)
Finally, the global model parameters are then aggregated by the re-weighted average, which ensures the contributions from different clients are proportional to their combined confidence scores and original weights. Let $\Theta_{\mathcal{H}}$ be the set of models from the honest clients, then the updated global model $\theta_{\text{global}}$ is obtained as:
$\theta_{\text{global}}=\frac{\sum_{i\in\mathcal{H}}w_{\text{final},i}\cdot\theta_{i}\cdot\ell_{i}}{\sum_{i\in\mathcal{H}}w_{\text{final},i}\cdot\ell_{i}}$ (14)
By interactively conducting the detection and selected aggregation, our proposed CAD framework effectively identifies and aggregates updates from honest clients, and ensure that the aggregation process gives more weight to clients with higher confidence scores, thus improving the overall performance and security of the FL model, and ultimately enhancing the robustness and reliability of the FL system.
Experiments
Experimental Setup
Datasets
We evaluated CAD model with CIFAR-10 (Krizhevsky and Hinton 2009), MNIST (LeCun et al. 1998), and Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017) benchmarks. which are common in prior studies. We employed a heterogeneous partitioning strategy to create non-IID local data where samples and class distributions are unbalanced among clients. The factor is sampled from a Dirichlet distribution $\texttt{Dir}_{N}(\alpha)$, with $\alpha$ determining the level of heterogeneity, where higher $\alpha$ values reduce heterogeneity. We set $\alpha=0.1$ to aligned with previous works.
DNN Models
We aligned our models with those employed in established previous works, namely three representative DNN models: AlexNet (Krizhevsky, Sutskever, and Hinton 2012), VGG-11 (Simonyan and Zisserman 2015), and a Multi-Layer Perceptron (MLP) (Popescu et al. 2009) with numer of layer $\{784,512,10\}$. Specifically, AlexNet and VGG-11 were used as the global model architectures for CIFAR-10, FC for MNIST, and AlexNet for Fashion-MNIST.
Table 2: Robustness to various intensities of attack on CIFAR-10 with VGG-11 Model. Accuracy of different defense methods under mild (25%), moderate (50%), and severe (75%) attack intensities. Best results are shown in bold.
Intensity Defence Label Shuffle LIE Min-Max Min-Sum
Full Partial Full Partial Full Partial Full Partial
Mild (25%) FedAvg 68.32 68.60 49.84 52.89 55.40 64.37 49.44 59.20
MKrum 66.84 66.25 29.94 28.12 44.59 36.90 57.16 43.17
TrimMean 65.20 65.20 31.63 34.22 53.69 61.35 51.58 60.33
AFA 68.88 69.09 69.09 68.80 68.89 68.71 68.99 69.17
FLDetector 39.74 46.32 61.94 56.92 32.66 32.37 25.93 35.61
DeFL 67.95 68.57 48.02 51.51 65.67 69.04 62.97 68.74
CAD (Ours) 69.36 69.47 69.67 69.18 69.85 69.43 69.94 69.43
Moderate (50%) FedAvg 62.91 62.70 12.94 15.04 19.02 18.37 21.40 22.25
MKrum 10.40 11.99 10.15 10.00 11.00 13.93 17.19 17.15
TrimMean 37.26 36.88 11.55 12.35 16.92 17.27 17.49 17.20
AFA 60.32 60.85 10.42 10.56 16.65 17.89 15.42 17.42
FLDetector 34.84 38.81 14.41 - 23.61 - 20.47 -
DeFL 63.02 62.31 23.17 23.95 21.60 - 48.26 20.76
CAD (Ours) 64.27 64.19 64.44 64.96 64.00 65.07 64.75 65.13
Severe (75%) FedAvg 50.51 53.46 10.56 12.08 16.71 17.72 17.79 17.60
MKrum 11.92 10.80 10.03 10.60 13.34 13.84 15.84 15.67
TrimMean 21.17 20.23 10.01 10.42 14.57 15.03 17.52 16.60
AFA 20.23 22.17 10.06 10.51 11.54 11.62 13.84 10.03
FLDetector 18.33 18.04 11.04 12.08 16.71 17.72 17.79 -
DeFL 52.18 52.72 16.76 - 20.11 - 19.42 20.02
CAD (Ours) 57.68 57.79 56.75 56.54 56.96 57.04 56.88 56.48
Baseline Defenses
We compared CAD with state-of-the-art defenses including: FedAvg (McMahan et al. 2017), Mkrum (Blanchard et al. 2017), TrimMean (Yin et al. 2018), AFA (Muñoz-González, Co, and Lupu 2019), FLDetector (Zhang and et al. 2022), DeFL (Yan et al. 2023). We consider two settings regarding the adversary’s knowledge as prior works: (a) Full, where the adversary knows the gradients of benign clients, and (b) Partial, where the adversary is unaware of the gradient updates shared by benign clients.
Malicious Client Attacks
We evaluated CAD model on both model poisoning attacks and data poisoning attacks:
• Label Shuffling (LS): Malicious clients randomly reassign labels of their local data, creating noisy and misleading training data that reduces the global model’s accuracy.
• Little is Enough (LIE): Adds small amounts of noise to the average of benign gradients. The adversary computes the average $\mu$ and standard deviation $\sigma$ of benign gradients, then applies a coefficient $z$ to obtain the malicious update $\mu+z\sigma$.
• Min-Max (MM): Computes malicious gradients such that their maximum distance from any gradient is upper bounded by the maximum distance between any two benign gradients, placing them close to the benign gradient cluster.
• Min-Sum (MS): Ensures the sum of squared distances of malicious gradients from all benign gradients is upper bounded by the sum of squared distances among benign gradients. Malicious gradients are kept identical to maximize impact.
Implementation Details
We employed a total of $N=50$ clients in our experiments. Each client was trained for 20 epochs using stochastic gradient descent to update its local model. The number of training rounds was set to 200 to ensure convergence for all DNN models on all datasets considered in this paper. The learning rate was initialized at 0.01 and decayed by a constant factor of 0.99 after each communication round. The batch size was set to 16, with a weight decay of $1\times 10^{-3}$. The percentage of malicious clients was set to three levels: mild (10%), moderate (25%), and severe (75%), to illustrate the robustness of our CAD model. For data poisoning attacks, confidence scores were directly obtained from the model’s output probabilities on the training data. For model poisoning attacks, confidence scores were obtained by inferencing on local training data before sending it to the parameter server. We implemented our defenses, in PyTorch on a single NVIDIA RTX A100 GPU.
Experimental Results
Comparison on Defense Robustness under Different Attack Intensities
We evaluated the robustness of our CAD model and other state-of-the-art defenses under different levels of attack intensity with the CIFAR-10 dataset and the VGG model. The results, as shown in Table 2, indicate that under mild attacks (25% malicious clients), most defense methods perform reasonably well, with our CAD model achieving the highest accuracy across all attack types. However, as the intensity of the attacks increases to moderate (50%) and severe (75%), the performance of most methods significantly deteriorates. In contrast, our CAD model consistently maintains high accuracy and robustness by outperforming other methods under all attack intensities. This demonstrates the effectiveness of our confidence-based detection and re-weighted aggregation approach in mitigating the impact of various attack types, even under severe attack conditions.
Defense Robustness across Various Models and Datasets
Table 3 compares the performance of CAD with other state-of-the-art defense methods across different models and datasets. The results demonstrate that CAD consistently outperforms other defenses with high accuracy under various attacks. For MNIST with MLP, most defense methods perform well under label shuffling, but CAD achieves the highest accuracy across all attack types. Similarly, for Fashion MNIST with AlexNet, CAD maintains superior performance, especially against LIE and MM attacks, where other methods significantly collapsed in accuracy. In the CIFAR-10 dataset, both with VGG-11 and AlexNet models, the robustness of CAD stands out. While other defenses struggle to converge, particularly under LIE and MM attacks, CAD consistently achieves the best results. This highlights the effectiveness of our method in providing robust defense against various attacks, even under different model and dataset combinations. These results demonstrated that CAD is not only effective in mitigating different types of attacks but also adaptable to various models and datasets, thereby making it a reliable defense mechanism in FL systems.
Table 3: Defense robustness across various models and datasets. Accuracy of different defense methods under various attacks across different models and datasets. Best results are shown in bold.
Model Defence LS LIE MM MS
MNIST (MLP) FedAvg 97.36 52.45 95.81 94.47
Mkrum 11.46 10.55 92.64 71.07
TrimMean 92.86 15.49 93.07 82.15
AFA 96.73 11.35 94.33 92.70
FLDetector 96.92 45.60 93.24 92.36
DeFL 97.47 88.90 94.29 93.29
CAD (Ours) 97.83 97.08 97.87 97.73
Fashion MNIST (AlexNet) FedAvg 88.06 15.66 78.71 74.58
Mkrum 10.02 10.00 27.13 18.61
TrimMean 75.17 10.00 56.02 53.31
AFA 88.56 10.00 23.70 23.70
FLDetector 76.51 10.00 78.37 74.98
DeFL 88.25 10.11 87.73 86.83
CAD (Ours) 88.79 87.86 88.43 88.32
Cifar10 (VGG-11) FedAvg 62.91 12.94 19.02 21.40
Mkrum 10.40 10.15 11.00 17.19
TrimMean 37.26 11.55 16.92 17.49
AFA 60.32 10.42 16.65 15.42
FLDetector 34.84 14.41 23.61 20.47
DeFL 63.02 23.17 21.60 48.26
CAD (Ours) 64.27 64.44 64.00 64.75
Cifar10 (AlexNet) FedAvg 59.63 10.16 19.86 19.75
Mkrum 11.77 10.00 17.61 16.83
TrimMean 33.80 10.00 19.34 17.44
AFA 57.02 10.01 18.68 19.47
FLDetector 32.47 10.00 17.24 17.17
DeFL 59.42 18.98 30.69 49.05
CAD (Ours) 62.76 62.90 62.92 62.60
Conclusion
In this paper, we proposed a Confidence-Aware Defense (CAD) model, a novel defense mechanism for federated learning that achieve malicious detection to address both model poisoning and data poisoning attacks. The proposed CAD model estimates per-client confidence scores to reflect prediction uncertainty, which are then referred to identify clients into honest and malicious categories. By re-weighting the contributions of honest clients based on their confidence scores, CAD model enhances the robustness and accuracy of the global model. More importantly, CAD can effectively detect and mitigate various attack conditions by focusing on client confidence scores. Comprehensive experiments on multiple datasets and models demonstrate that CAD significantly improves the accuracy, stability, and robustness of federated learning systems compared to existing state-of-the-art methods.