### v2

# Mitigating Malicious Attacks in Federated Learning via Confidence-aware Defense

Abstract
Federated Learning (FL) is an distributed machine learning diagram that enables multiple clients to collaboratively train a global model without sharing their private local data. However, FL systems are vulnerable to attacks that are happening in malicious clients through data poisoning and model poisoning who can deteriorate the aggregated global model’s performance. Existing defense methods typically focus on mitigating specific types of poisoning, and are often ineffective against unseen types of attack. These methods also assume an attack happened moderately while is not always holds true in real. Consequently, these methods can significantly fail in terms of accuracy and robustness when detecting and addressing updates from attacked malicious clients. To overcome these challenges, in this work, we propose a simple yet effective framework to detect malicious clients, namely Confidence-Aware Defense (CAD), that utilizes the confidence scores of local models as criteria to evaluate the reliability of local updates. Our key insight is that malicious attacks, regardless of attack type, will cause the model to deviate from its previous state, thus leading to increased uncertainty when making predictions. Therefore, CAD is comprehensively effective for both model poisoning and data poisoning attacks by accurately identifying and mitigating potential malicious updates, even under varying degrees of attacks and data heterogeneity. Experimental results demonstrate that our method significantly enhances the robustness of FL systems against various types of attacks across various scenarios by achieving higher model accuracy and stability.
Introduction
Federated Learning (FL) (Zhang et al. 2021) is a distributed machine learning paradigm that allows multiple clients to collaboratively train a global model without sharing their local data. FL offers significant privacy protection, therefore making it highly attractive for privacy-sensitive applications (Rodríguez-Barroso et al. 2023). However, the decentralized nature of FL also introduces security challenges: malicious attacks happened in distributed local clients can degrade the global model’s performance through data poisoning and model poisoning attacks, leading to degraded predictions and potentially catastrophic consequences (Fang and Chen 2023; Kumar, Mohan, and Cenkeramaddi 2023). As illustrated in Fig 1, malicious attacks can significantly hinder the convergence of DNN model and even completely cause collapse. Therefore, developing effective methods to detect and defend against malicious attacks is crucial in ensuring the security and robustness of FL systems.
Figure 1: Impact of malicious attacks on model convergence under the “Little Is Enough” attack at various intensities. The FedAvg aggregation strategy degrades significantly, while our CAD consistently maintains high accuracy and robustness.
There are two main types of attacks in a FL system: model poisoning and data poisoning, categorized by where the attack occurs. Model poisoning attacks involve directly manipulating the model updates sent by the clients to the central server. Among these, Byzantine attacks (Shi et al. 2022) are the most representative, where malicious clients intentionally send arbitrary or random updates to disrupt the learning process and significantly degrade the performance of the global model (Fang et al. 2020; Bhagoji et al. 2019). In contrast, data poisoning attacks occur when malicious clients intentionally corrupt their local training data to mislead the global model during the training process, such as label shuffle, where the data-label pairs are randomly reassigned that can confuse the globally updated local models (Tolpegin et al. 2020; Xie et al. 2020). Both attack types pose significant threats, and cause the global model to make incorrect predictions, degrading its accuracy and robustness. These challenges underscore the urgent need for effective detection and defense mechanisms capable of addressing various attack types.
Existing defense mechanisms primarily target specific types of poisoning attacks, under a moderate attack. Byzantine-robust methods such as Krum (Blanchard et al. 2017) and Trimmed Mean (Yin et al. 2018) aim to mitigate the impact of outliers among client updates but often fall short when dealing with data poisoning attacks. However, these methods generally assume that the majority of clients are honest, which may not hold true in practical scenarios where the proportion of malicious clients is significant (Fang et al. 2020). Recent advanced methods are primarily driven by the analysis of fine-grained model consistency. In that context, FLDetector (Zhang and et al. 2022) improves upon existing defenses by detecting malicious clients through the consistency of their model updates across historical iterations, and flags clients whose updates deviate significantly. DeFL (Yan et al. 2023) focuses on critical learning periods within the training process to determine the gradient changes and remove malicious clients. FedRoLA (Gang Yan and Li 2024) enhances robustness by analyzing the per-layer consistency of model updates and conducts aggregation based on the alignment of updates at each layer. However, these methods overlook the uncertainty in model updates and fail to address various types of both poisoning attacks comprehensively.
Given the inherent uncertainty of FL systems and the attacks’ complexity, being capable of addressing only certain types of attack is far from sufficient. A robust and comprehensive defense approach is needed to ensure the integrity and reliability of the global model across all scenarios, including various types of attacks with differing levels of severity and degrees of data heterogeneity. To meet this demand, we propose a simple yet effective defense model, termed Confidence-Aware Defense (CAD), based on the confidence scores of clients’ local models, which measures how certain a model is about its predictions, with higher scores indicating greater certainty and lower scores reflect higher uncertainty. Our key insight is that malicious attacks, regardless of attack type, will cause the model to deviate from its previous state, thus leading to increased uncertainty in its predictions. Therefore, CAD is comprehensively effective for both model poisoning and data poisoning attacks. Specifically, the proposed approach includes the following steps: 1. Collecting confidence scores of each client update: During each training round, we collect the confidence scores of model updates from each client. 2. Establishing confidence boundaries: Based on the collected confidence scores, we assess the uncertainty of each client model update and set the boundaries as a reference. 3. Detecting and handling malicious updates: Based on the boundaries, we identify updates with high uncertainty and take appropriate actions, such as discarding or adjusting these updates.
Our key contributions are summarized as follows:
  • We introduce a simple yet effective confidence-based malicious client detection method to identify if any attack occurs in a local client. Our approach is versatile and applicable to both model poisoning and data poisoning attacks, under various data heterogeneities.
  • By detecting malicious attacks on client model updates through confidence scores, our method significantly improves the robustness of FL systems against attacks of different intensities. Whether facing moderate (25% malicious clients), severe (50% malicious clients), or extreme (75% malicious clients) activities, our approach ensures higher model accuracy and stability.
  • We validate the proposed method through extensive experiments on multiple datasets and demonstrate its effectiveness in enhancing the security and performance of FL models against a wide range of attacks. The results show that our confidence-based defense mechanism outperforms existing methods in terms of robustness and accuracy, in detecting and mitigating malicious updates.
Related Work
Federated Learning (FL)
FL is a decentralized machine learning diagram that enables multiple clients to collaboratively train a shared global model while keeping their local data private. In FL, the learning process is divided into two main stages: local training and global aggregation. During the local training phase, each client trains a local model using its private dataset. This involves updating the model parameters based on the local data, which protects privacy by not allowing raw data to leave the client’s device. Once the local training is complete, the clients send their updated model parameters to a central server for global aggregation, in which the central server aggregates the received model updates from participating clients to form a new global model. This aggregation can be done by various approaches, the most common is the Federated Averaging (FedAvg) algorithm (McMahan et al. 2017). The FedAvg algorithm computes the weighted average of the model parameters from all clients to form a global model. Despite the advantages of FL in preserving data privacy, the decentralized nature of this learning paradigm makes it vulnerable to various types of attacks either caused by data or model poisoning. Therefore, it is crucial to develop efficient and robust defense mechanisms that can identify and mitigate the impact of malicious activities.
Malicious Client Attacks
FL models are susceptible to various types of attacks from malicious clients, which can be broadly categorized into data poisoning and model poisoning.
Data Poisoning Attacks: In data poisoning attacks, malicious clients intentionally manipulate their local training data to degrade the performance of the global model. Common strategies include label flipping, where the labels of training samples are deliberately swapped, and label shuffling, where the labels of training samples are randomly reassigned. These attacks aim to bias the model’s training process, and further lead to inaccurate or harmful predictions. Rahman et al. Rahman et al. (2023) highlights how local model poisoning through data manipulation can significantly impact the robustness of federated learning systems. Similarly, Smith et al. (Smith and Brown 2023) examine clean-label poisoning attacks, where the manipulated data appears legitimate but is designed to subvert the performance of the model.
Model Poisoning Attacks: In contrast to data poisoning attacks, model poisoning attacks refer to malicious clients directly sending manipulated model updates to the central server during the global aggregation phase. These attacks are conducted by altering the model parameters in a way that maximizes the negative impact on the global model. Zhang et al. (Zhang and Chen 2023) have demonstrated how backdoor attacks could be performed by injecting specific patterns into the model updates, enabling attackers to trigger incorrect behaviors under certain conditions. Lee et al. (Lee and Park 2023) analyze the adversarial impact of model attacks on federated learning models. Additionally, Byzantine attacks, which involve arbitrary or adversarial behavior by clients, can also be classified under model poisoning when clients send random or adversarial updates to the central server. Recent Byzantine-robust methods like Krum++ (Wang and Zhou 2023) and Enhanced Trimmed Mean (Liu and Zhao 2023) aim to filter out these abnormal updates but can struggle when the attacks are sophisticated.
Defense Against Malicious Attacks
To detect and mitigate the aforementioned attacks, numerous works have been proposed recently (Fang and Chen 2023; Yan et al. 2023; Zhang and et al. 2022; Gang Yan and Li 2024). DeFL (Yan et al. 2023) is proposed as a defense mechanism that secures critical learning periods during the training process to reduce the impact of model poisoning attacks. Zhang et al. (Zhang and et al. 2022) introduced FLDetector to integrate statistical analysis, anomaly detection, and machine learning methods to enhance the accuracy of detecting poisoned updates. FedRoLA (Gang Yan and Li 2024) strengthens robustness by analyzing per-layer consistency in model updates and conducting aggregation based on the alignment of updates at each layer. Despite the improved accuracy in detecting malicious clients, these existing methods are often tailored to specific types of attacks. Additionally, they typically rely on the fragile assumption that the attack intensity is moderate, meaning less than 25% of the participating clients are malicious, which may not hold in a practical FL system where higher proportions of malicious clients can be present. To resolve these limitations, in this paper, we propose the CAD model to address both types of attacks and varying intensities.
Model Confidence and Uncertainty
Confidence scores (Gass and Joel 1981) measure a model’s certainty regarding its predictions. Higher confidence scores indicate greater confidence, while lower scores suggest higher uncertainty. It provides insights into the model’s decision-making process by identifying the predictions that the model is less sure about (Gal and Ghahramani 2016; Lakshminarayanan, Pritzel, and Blundell 2017).
Super Loss is a confidence-aware loss designed for curriculum learning, which incorporates confidence scores to determine the reliability of each training instance (Castells, Weinzaepfel, and Revaud 2020). it introduces a regularization to adjust the loss dynamically based on the confidence score (${\sigma}^{j}$) for sample $j$ in the training set $\mathcal{D}$. The Super Loss is defined as:
$\mathcal{L}_{\text{super}}=\frac{1}{N}\sum_{j=1}^{N}\left(\frac{1}{{\sigma^{j}}^{2}}\mathcal{L}_{j}+\log\sigma_{j}\right),$ (1)
where $\mathcal{L}_{i}$ is the task loss for sample $i$, and ${\sigma}^{j}$ is the confidence score, and $N$ is the number of samples. The confidence score ${\sigma}^{j}$ is obtained by the following closed-form solution:
${\sigma}^{j}=\exp\left(-\text{W}\left(\frac{1}{2}\max\left(-\frac{2}{e},\frac{\mathcal{L}_{i}-\log(C)}{\lambda}\right)\right)\right),$ (2)
where $\log(C)$ is the constant, and $C$ is set to the number of classes, and $\lambda$ is the regularization parameter. In this work, motivated by the observation that malicious attack will disrupt the local model’s convergence against the original global initialization, and making the learned model less confident in the output, we propose to use per-client confidence accumulation as the index to determine the quality of the learned model. Note that we use Eq. (2) solely for estimating the confidence score, in contrast to the original design of Super Loss which uses it as a regularization term. The rationale is that our goal is to make the CAD model versatile and applicable to various tasks, and to explicitly demonstrate that the effectiveness stems from the detection algorithm rather than modifications to the loss function.
Methodology
Figure 2: The proposed Confidence-Aware Defense (CAD) framework. It encompasses five key steps: 1. Initialize the global model and distribute it to clients. 2. Clients train their personalized, confidence-aware local models. 3. Clients upload their local models and associated confidence scores to the server. 4. The server identifies potentially malicious clients through confidence clustering. 5. The server aggregates the local models of honest clients by re-weighted aggregation.
Problem Formulation
We consider an FL system with $n$ clients, each with its local dataset $\mathcal{D}_{i}$. The goal is to train a global model $\theta$ by aggregating the local updates $\theta_{i}$ from client $\{C_{i}\}_{i=1}^{n}$ while mitigating the impact caused by potential malicious clients, which can perform two types of attacks: Data poisoning attacks: These include label manipulation where the local dataset is intentionally corrupted to mislead the global model. Model poisoning attacks: Here, the model updates sent by the clients are manipulated to degrade the performance of the global model. To achieve this goal, we propose a Confidence-Aware Defense (CAD) mechanism based on the per-client confidence scores ($\{\sigma_{i}\}_{i=1}^{n}$), which reflects the uncertainty of the local model’s predictions and helps in identifying malicious updates. The notations used in formulating CAD are summarized in Table 1.
Table 1: Table of notations used in this paper.
Notation Description
$n$ Number of clients
$\mathcal{D}_{i}$ Local dataset of client $i$
$\theta$ Global model parameters
$\theta_{i}$ Local model parameters of client $i$
$\mathcal{L}_{i}^{j}$ Individual loss for sample $j$ in $\mathcal{D}_{i}$
$\sigma_{i}$ Confidence score for client $i$
$S$ Set of confidence scores from all clients
$\mu_{\text{lower}},\mu_{\text{upper}}$ Centroids from k-means clustering
$\mathcal{H}$ Set of honest clients
$\mathcal{M}$ Set of malicious clients
$\mathcal{L}_{\mathcal{H}}$ Data lengths of honest clients
$W_{\text{orig}}$ Original weights based on data lengths
$R$ Re-weight regularization factors
$r_{\text{norm},i}$ Normalized regularization factor
$w_{\text{final},i}$ Final weight for client $i$
W Lambert W function
Confidence-Driven Detection
Client Confidence Estimation.
For each client, the confidence score is estimated with their local dataset $\mathcal{D}_{i}$. Referring to Eq. (2), for each local client, we calculate the accumulated confidence score for all samples in the local dataset and then compute an average to obtain the per-client confidence score $\sigma_{i}$. by averaging the accumulated confidence scores for client $C_{i}$, which is the average confidence scores among all samples in its dataset:
$\sigma_{i}=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\sigma^{j}_{i},$ (3)
where $N_{i}$ is the number of samples in $\mathcal{D}_{i}$. This $\sigma_{i}$ reflects the overall uncertainty of the local model’s predictions for client $i$. Once the confidence scores for all clients have been computed, they will be normalized to ensure comparability across different clients by min-max scaling, which transforms the confidence scores into a uniform range [0, 1]. Let $S=\{\sigma_{i}\}_{i=1}^{n}$ represent the set of confidence scores from $n$ clients. The normalized confidence scores $S_{\text{norm}}$ are formulated as:
$S_{\text{norm}}=\frac{S-\min(S)}{\max(S)-\min(S)}.$ (4)
Clustering Based Detection.
After normalizing the per-client confidence scores, the next step is to cluster the clients into groups of honest and potentially malicious clients. The rationale for such clustering is based on the assumption that clients can be divided into two groups: honest clients with higher confidence scores and potentially malicious clients with lower confidence scores. To achieve this, we apply two-way k-means clustering to the normalized confidence scores $S_{\text{norm}}$. By setting the number of clusters to 2, we obtain two centroids to establish the lower bound and upper bound:
$\{\mu_{\text{lower}},\mu_{\text{upper}}\}=\text{k-means}(S_{\text{norm}},2).$ (5)
Once the centroids are obtained, the following step is to classify the clients as honest or malicious based on the distance in terms of the normalized confidence score to the cluster centroids. Specifically, a client $i$ is deemed honest if its normalized confidence score $\sigma_{i}$ is closer to the centroid $\mu_{\text{upper}}$ (representing higher confidence) than to the centroid $\mu_{\text{lower}}$ (representing lower confidence). Conversely, a client is classified as malicious if the opposite holds true:
$\mathcal{H}$ =$=\{i\mid|\sigma_{i}-\mu_{\text{upper}}|<|\sigma_{i}-\mu_{\text{lower}}|\},$ (6)
$\mathcal{M}$ =$=\{i\mid|\sigma_{i}-\mu_{\text{lower}}|\leq|\sigma_{i}-\mu_{\text{upper}}|\}.$
By classifying clients in this manner, we can focus on aggregating updates from only honest clients and get rid of malicious ones, thereby enhancing the robustness and security of an FL system.
Re-weighted Aggregation
Based on the identified set of malicious clients, the next step is to perform weight aggregation. Let $\Theta_{\mathcal{H}}$ denote the set of honest local model updates, and $\mathcal{L}_{\mathcal{H}}$ represent the corresponding set of data lengths for these honest clients. These are defined as:
$\Theta_{\mathcal{H}}=\{\theta_{i}\mid i\in\mathcal{H}\},$ (7)
=$\mathcal{L}_{\mathcal{H}}=\{\ell_{i}\mid i\in\mathcal{H}\},$ (8)
by which we isolate the local updates of clients that are considered trustworthy, so as to ensure that only reliable information will be used in the aggregation process. Subsequently, we define the original weights $w_{\text{orig},i}$ based on the data lengths of the honest clients. The original weight for each client is calculated as the ratio of the client’s data length to the total data length of all honest clients:
$W_{\text{orig}}=\{w_{\text{orig},i}\}_{i\in\mathcal{H}}=\{\frac{\ell_{i}}{\sum_{j\in\mathcal{H}}\ell_{j}}\}_{i\in\mathcal{H}}.$ (9)
In addition, to adaptively optimize the weight based on the trustworthiness (confidence) of each honest client, we apply a re-weighting strategy, for which a re-weighting regularization factor $r_{i}$ is then calculated for each honest client based on their normalized confidence scores. The set of re-weighting regularization factors for honest clients is denoted as:
$R=\{\sigma_{i}\mid i\in\mathcal{H}\}.$ (10)
The factors are then normalized:
$r_{\text{norm},i}=\frac{\sigma_{i}}{\sum_{j\in\mathcal{H}}\sigma_{j}},$ (11)
so that one can ensure that the aggregation process give appropriate importance to individual client models. Once calculated the original weights and the re-weighting regularization factors, we combine both to obtain the final weights as:
$w_{\text{final},i}=r_{\text{norm},i}\cdot w_{\text{orig},i}.$ (12)
As a final step, the global model parameters are then aggregated by the re-weighted average, which ensures the contributions from different clients are proportional to their combined confidence scores and original weights. Let $\Theta_{\mathcal{H}}$ be the set of models from the honest clients, then the updated global model $\theta_{\text{global}}$ is obtained as:
$\theta_{\text{global}}=\frac{\sum_{i\in\mathcal{H}}w_{\text{final},i}\cdot\theta_{i}\cdot\ell_{i}}{\sum_{i\in\mathcal{H}}w_{\text{final},i}\cdot\ell_{i}}.$ (13)
By interactively conducting this detection and selected aggregation, our proposed CAD model effectively identifies and aggregates updates from honest clients, while ensuring that the aggregation process gives more weight to clients with higher confidence scores, thus improving the overall performance and security of the FL model.
Experiments
Experimental Setup
Table 2: Robustness to various intensities of attack on CIFAR-10 with VGG-11 model. Best results are shown in bold.
Intensity Defence Min-Max Min-Sum LIE Label Shuffle
Full Partial Full Partial Full Partial Full Partial
Moderate (25%) FedAvg 55.40 64.37 49.44 59.20 49.84 52.89 68.32 68.60
TrimMean 53.69 61.35 51.58 60.33 31.63 34.22 65.20 65.20
AFA 68.89 68.71 68.99 69.17 69.09 68.80 68.88 69.09
FLDetector 32.66 32.37 25.93 35.61 61.94 56.92 39.74 46.32
DeFL 65.67 69.04 62.97 68.74 48.02 51.51 67.95 68.57
FedRoLA 69.23 69.03 68.59 68.76 68.67 68.95 68.48 68.59
CAD (Ours) 69.85 69.43 69.94 69.43 69.67 69.18 69.36 69.47
Severe (50%) FedAvg 19.02 18.37 21.40 22.25 12.94 15.04 62.91 62.70
TrimMean 16.92 17.27 17.49 17.20 11.55 12.35 37.26 36.88
AFA 16.65 17.89 15.42 17.42 10.42 10.56 60.32 60.85
FLDetector 23.61 20.02 20.47 20.73 14.41 11.96 34.84 38.81
DeFL 21.60 27.50 48.26 48.26 23.17 23.95 63.02 62.31
FedRoLA 19.52 19.60 21.40 22.10 13.42 14.17 62.52 61.98
CAD (Ours) 64.00 65.07 64.75 65.13 64.44 64.96 64.27 64.19
Extreme (75%) FedAvg 16.71 17.72 17.79 17.60 10.56 12.08 50.51 53.46
TrimMean 14.57 15.03 17.52 16.60 10.01 10.42 21.17 20.23
AFA 11.54 11.62 13.84 10.03 10.06 10.51 20.23 22.17
FLDetector 16.71 17.72 17.79 16.31 11.04 12.08 18.33 18.04
DeFL 20.11 19.44 19.42 20.02 16.06 16.76 52.18 52.72
FedRoLA 16.71 16.86 17.79 17.59 10.00 10.00 54.39 52.46
CAD (Ours) 56.96 57.04 56.88 56.48 56.75 56.54 57.68 57.79
Datasets.
We evaluated CAD model a wide range of benchmarks including: CIFAR-10 (Krizhevsky and Hinton 2009), MNIST (LeCun et al. 1998), and Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017) datasets, which are common in prior studies (Yan et al. 2023; Gang Yan and Li 2024). We employed a heterogeneous partitioning strategy to create non-IID local data where samples and class distributions are unbalanced among clients. The factor is sampled from a Dirichlet distribution $\texttt{Dir}_{N}(\alpha)$, with $\alpha$ determining the level of heterogeneity, where a smaller $\alpha$ value refers to high heterogeneity (Shang et al. 2022). We set $\alpha=0.5$ by default in the following experiments. We evaluated the performance by classification accuracy.
DNN Models.
We tested CAD model on three representative DNN models including: AlexNet (Krizhevsky, Sutskever, and Hinton 2012), VGG-11 (Simonyan and Zisserman 2015), and a Multi-Layer Perceptron (MLP) (Popescu et al. 2009) with layers consisting of 784, 512, and 10 neurons respectively. Specifically, AlexNet and VGG-11 were employed for the CIFAR-10 dataset, the MLP for MNIST, and AlexNet for Fashion-MNIST.
Baseline Defenses.
We compared CAD with state-of-the-art defense models including: TrimMean (Yin et al. 2018), AFA (Muñoz-González, Co, and Lupu 2019), FLDetector (Zhang and et al. 2022), DeFL (Yan et al. 2023), FedRoLA (Gang Yan and Li 2024). Additionally, we included FedAvg (McMahan et al. 2017) as a baseline method for comparison. We consider two settings regarding the adversary’s knowledge (Yan et al. 2023): (a) Full, where the adversary knows the gradients of honest clients, and (b) Partial, where the adversary is unaware of the gradient updates shared by honest clients. We evaluated the detection performance with True Positive Rate (TPR) and False Positive Rate (FPR).
Malicious Client Attacks.
We tested CAD model on both model poisoning attacks and data poisoning attacks:
• Label Shuffling (LS): Malicious clients randomly reassign labels of their local data, creating noisy and misleading training data that reduces the global model’s accuracy.
• Little is Enough (LIE): Adds small amounts of noise to the average of honest gradients. The adversary computes the average $\mu$ and standard deviation $\sigma$ of honest gradients, then applies a coefficient $z$ to obtain the malicious update $\mu+z\sigma$.
• Min-Max (MM): Computes malicious gradients such that their maximum distance from any gradient is upper bounded by the maximum distance between any two honest gradients, placing them close to the honest gradient cluster.
• Min-Sum (MS): The sum of squared distances of malicious gradients from all honest gradients is upper bounded by the sum of squared distances among honest gradients. Malicious gradients are kept identical to maximize impact.
Implementation Details.
We simulated $N=50$ clients in the experiments. Each client was trained for 20 epochs with stochastic gradient descent to update its local model. The number of training rounds was set to 200 to ensure convergence for all DNN models. The learning rate was set to 0.01. The batch size was set to 16, with a weight decay of $1\times 10^{-3}$. We evaluated the robustness of our CAD model under three degrees of attack severity defined by the percentage of malicious clients: moderate (25%), severe (50%), and extreme (75%). For data poisoning attacks, confidence scores were directly obtained during model training. For model poisoning attacks, confidence scores were obtained by further running inference on local training data before sending it to the parameter server. We implemented our defenses, in PyTorch (Paszke et al. 2019) on a single NVIDIA A100 GPU.
Experimental Results
Comparison on Defense Robustness under Different Attack Intensities.
We evaluated the robustness of the proposed CAD model and other state-of-the-art defenses under different levels of attack intensity with the CIFAR-10 dataset with the VGG model. The results, as shown in Table 2, indicate that under moderate attacks (25% malicious clients), most defense methods perform reasonably well, with our CAD model achieving the highest accuracy across all attack types. However, as the intensity of the attacks increases to severe (50%) and extreme (75%), the performance of most methods significantly deteriorates. In contrast, the CAD model consistently maintains high accuracy and robustness by outperforming other methods under all attack intensities. This demonstrates the effectiveness of our confidence-based detection and re-weighted aggregation approach in mitigating the impact of various attack types, even under severe attack conditions.
Applicable to Various Models and Datasets.
Table 3 compares the performance of CAD with other state-of-the-art defense methods across different models and datasets. The results demonstrate that CAD consistently outperforms other defenses with high accuracy under various attacks. For MNIST with MLP, most defense methods perform well under label shuffling attacks, but CAD achieves the highest accuracy across all attack types. Similarly, for Fashion MNIST with AlexNet, CAD maintains superior performance, especially against LIE and MM attacks, where other methods significantly collapsed in accuracy. In the CIFAR-10 dataset, both with VGG-11 and AlexNet models, the robustness of CAD stands out. While other defenses struggle to converge, particularly under LIE and MM attacks, CAD consistently achieves the best results. This highlights the effectiveness of our method in providing robust defense against various attacks, even under different model and dataset combinations. These results demonstrated that CAD is not only effective in mitigating attacks for a specific benchmark but also adaptable to various models and datasets, thereby making it a reliable defense method.
Table 3: Defense robustness across various models and datasets. Accuracy of different defense methods with various attacks across different models and datasets under the severe attack setting. The best results are shown in bold.
Model Defence MM MS LIE LS
Cifar10 (VGG-11) FedAvg 19.02 21.40 12.94 62.91
AFA 16.65 15.42 10.42 60.32
FLDetector 23.61 20.47 14.41 34.84
DeFL 21.60 48.26 23.17 63.02
FedRoLA 19.52 21.40 13.42 62.52
CAD (Ours) 64.00 64.75 64.44 64.27
Cifar10 (AlexNet) FedAvg 19.86 19.75 10.16 59.63
AFA 18.68 19.47 10.01 57.02
FLDetector 17.24 17.17 10.00 32.47
DeFL 30.69 49.05 18.98 59.42
FedRoLA 19.52 19.75 10.16 59.33
CAD (Ours) 62.92 62.60 62.90 62.76
Fashion MNIST (AlexNet) FedAvg 78.71 74.58 15.66 88.06
AFA 23.70 23.70 10.00 88.56
FLDetector 78.37 74.98 10.00 76.51
DeFL 87.73 86.83 10.11 88.25
FedRoLA 78.09 74.58 15.66 88.05
CAD (Ours) 88.43 88.32 87.86 88.79
MNIST (MLP) FedAvg 95.81 94.47 52.45 97.36
AFA 94.33 92.70 11.35 96.73
FLDetector 93.24 92.36 45.60 96.92
DeFL 94.29 93.29 88.90 97.47
FedRoLA 95.81 94.47 52.45 97.40
CAD (Ours) 97.87 97.73 97.08 97.83
Robustness to Data Heterogeneity.
Figure 3: Robustness to data heterogeneity by varying degrees of Non-IID data, as controlled by the parameter $\alpha$ in the Dirichlet distribution $\texttt{Dir}(\alpha)$.
We further investigated the robustness of CAD across different levels of data heterogeneity by varying the key parameter $\alpha$ in the Dirichlet distribution $\text{Dir}(\alpha)$. A smaller $\alpha$ indicates a higher degree of non-IID data. We simulated a moderate LIE attack and a severe Label Shuffle attack on the Cifar10 dataset with VGG-11 model. As shown in Fig. 3, as the non-IID degree decreases (i.e., , as $\alpha$ increases), the global model accuracy improves at a much higher rate compared to FedAvg, as indicated by the blue shaded area, consistently under both Full and Partial scenarios. These results confirm the robustness of our CAD model to varying data heterogeneity.
Malicious Clients Detection Accuracy.
We evaluated the detection performance of our proposed CAD model under LIE and label shuffle attack scenarios considering both moderate and severe intensities. The evaluation was conducted on the Cifar10 dataset with the VGG-11 model at full attack setting. As shown in Fig. 4, our CAD model can accurately detect malicious clients, and converge to high TPR and low FPR after only a few training rounds. Therefore, it can exclude malicious clients and lead to a significant improvement in overall performance, as shown in the blue line over the orange line. Its advantage is more obvious when more clients are being attacked, with detection accuracy remaining high and stable even under severe attack conditions while the FedAvg was collapsed. This demonstrates that CAD is effective under varying intensities and thereby significantly enhances model performance compared to the FedAvg baseline.
Figure 4: Malicious client detection accuracy evaluated on the Cifar10 dataset with the VGG-11 model.
Impact of Re-weighting.
We ablated the impact of the re-weighting operation in Eq. (12) and the results are shown in Fig. 5. It can be observed that the re-weighting operation can consistently enhance performance, especially at higher attack rates. The improvement is particularly clear in “Severe” and “Extreme” conditions, with detecting attacks like LIE and MinSum benefiting the most. This effectiveness is likely due to the fact that only a few honest clients remain under high attack rates, therefore, it is crucial to identify and prioritize them. This further demonstrates that the confidence score is a reliable metric for assessing the quality of client models, thereby supporting the underlying assumption of our confidence-aware CAD model.
Figure 5: Effect of re-weighting on model accuracy across different degrees of attack.
Conclusion
In this paper, we proposed the CAD model, an attack defense mechanism for FL that achieves accurate malicious client detection to address both model poisoning and data poisoning attacks. The proposed CAD model estimates per-client confidence scores to reflect the model’s uncertainty, which are then referred to identify clients into honest and malicious categories. By re-weighting the contributions of honest clients based on their confidence scores, CAD model can make the best use of honest clients even under severe attack when there is only a handful of honest clients available, so that it can enhance the robustness of the global model. Extensive experiments conducted on various benchmarks with non-IID data demonstrated that CAD can notably enhance the accuracy, stability, and robustness of FL systems.